in a laboratory, only 40% and 50% alcohol solutions are available. for an experement, charmaine must use alcohol with a concentration of 45%. if 200L of the solution are needed, how manny liters of each solution should be mixed
Answer by richwmiller(17219)

x+y=200,
0.4*x+0.5*y=0.45*200
0.4*x+0.5*y=90
x=200-y
0.4*(200-y)+0.5*y=90
80-0.4y+0.5*y=90
0.1*y=10
y=100 at 50%
x=200-y
x=100 at 40%
check
0.4*100+0.5*100=0.45*200
40+50=90
90=90
ok